# Minimizing Human Intervention in Online Classification

### v1

@@ F.2.1 Real-World Datasets
Quora Question Groups (QQG): Quora Question Groups is derived from the Quora Question Pairs dataset (Wang et al.,, 2017). The original dataset comprises over 400,000 question pairs, each annotated with a binary label indicating whether the questions are paraphrases of each other. To transform this pairwise paraphrase data into coherent question groups, we represented the dataset as a graph where nodes correspond to questions and edges indicate paraphrase relationships.
@@
ComQA [^6] https://paperswithcode.com/dataset/comqa (Abujabal et al.,, 2019): ComQA consists of 11,214 English questions collected from the WikiAnswers forum and grouped into 4,834 paraphrase clusters by crowd workers. By design, ComQA emphasizes reasoning-intensive phenomena—namely compositional queries, explicit temporal constraints, and comparative or superlative constructions—making it a rigorous benchmark for research on complex factoid question answering. The middle panel of Fig. 7 presents the group size distribution of this dataset.
@@
CQADupStack  [^7] https://github.com/D1Doris/CQADupStack (Hoogeveen et al.,, 2015): CQADupStack is a public benchmark available to researchers working on question-answering learning tasks. It consists of complete threads from twelve StackExchange subforums (android, english, gis, mathematica, physics, programmers, statistics, superuser, tex, unix, webmasters, wordpress) annotated with community‐marked duplicate links. To evaluate our classification approach across distinct expert domains, we apply it to the Mathematica, Physics, and Statistics parts of the dataset. The right panel of Fig. 7 presents the group size distribution of this dataset. The subset of the dataset we use comprises 99,785 questions organized into 74,519 groups. CQADupStack is the most challenging dataset used in our experiments, because of its relatively small average group size.
F.2.2 Large Language Models
@@
In this section, we describe the Large Language Models (LLMs) selected for our experiments, along with the way they were trained and fine-tuned. We chose three models that are widely recognized in the literature and frequently cited due to their strong performance across a range of Natural Language Processing tasks.
@@
E5 (Wang et al.,, 2022): EmbEddings from bidirEctional Encoder rEpresentations (E5) is available in three versions. The first version is initialized with MiniLM (Wang et al.,, 2020), the second is initialized with BERT-base (Reimers and Gurevych,, 2019), and the third is initialized with BERT-large. This model follows a bi-encoder architecture, where both the query and passage encoders are initialized with BERT. The training process consists of two stages. The first stage, called “unsupervised”, uses a large number of unlabeled data, including title and passage pairs from Wikipedia, questions and answers from Reddit, and more. The InfoNCE contrastive loss (van den Oord et al.,, 2018) is employed to minimize the distance between related queries and passages, while maximizing the distance between unrelated queries and passages. The second stage involves training the model on high-quality human-annotated data, such as NLI (Gao et al.,, 2021), MS-MARCO (Bajaj et al.,, 2016), and Natural Questions (Karpukhin et al.,, 2020). During this stage, the model is trained with a loss that combines the KL divergence between the probability distribution of the label, as given by a cross-encoder teacher model, and the probability distribution generated by our E5 student model, along with the InfoNCE contrastive loss. This second stage further improves model performance on benchmarks such as BEIR (Thakur et al.,, 2021) and MTEB (Muennighoff et al.,, 2022).
@@
NOMIC (Nussbaum et al.,, 2024): Nomic is initialized from BERT and modified to address long-context retrieval. Nomic consists of 100 million parameters and supports a sequence length of up to 2048. Nomic’s pretraining is divided into three stages. The first stage focuses on Masked Language Modeling to learn longer sequence representations. The subsequent stages are supervised and unsupervised, both employing the InfoNCE contrastive loss. This model was trained on a significantly larger dataset compared to the previous versions, encompassing both supervised and unsupervised phases. Nomic uses task-specific prefixes—such as search, search query document, classification, and clustering—to distinguish between the behaviors of each task. For the purpose of our work, we used the clustering prefix.
@@
Mistral_E5 (Wang et al.,, 2024): This is the first unidirectional decoder architecture we use for our work. The model is initialized from Mistral 7B Jiang et al., (2023) and consists of 7 billion parameters. The model takes as input the query $q^{+}$ and the task_definition and generates the instruction template:
@@
The query and document instruction templates are then passed to the LLM. We obtain the query and document embeddings, $h^{+}_{q}$ and $h^{+}_{d}$, from the last layer of the LLM at the [EOS] position. The model was trained on a large corpus of both original and synthetic data. The synthetic data was generated using advanced LLMs such as GPT-4.
F.2.3 Additional Real-World Experiments
In this section, we evaluate the performance of GHC (Algorithm 2), CC and a sequential $k$-means baseline using the datasets and LLMs described in Sections F.2.1 and F.2.2 respectively. The models we utilize are typically trained using a loss function based on cosine similarity. Since they output unit-normalized embeddings, maximizing cosine similarity is equivalent to minimizing Euclidean distance. Hence, the reliance of GHC on Euclidean distance naturally leverages the models’ cosine-trained embeddings. As explained in Section 5, we always initialize the algorithms by manually picking an example question from each group. For GHC, this skips the first phase where CHC is applied until every group is represented. We then run our algorithms on the remaining questions in the dataset considered.
To align with the Retrieval-Augmented Generation (RAG) literature (Lewis et al.,, 2020) and improve computational efficiency, it is sensible to consider the following modification of the GHC algorithm: instead of computing distances to the convex hulls $\hat{\mathcal{C}}_{i,t}$ of each set of expert queries $\mathcal{Q}_{i,t}$, we compute the distance to the nearest query in $\mathcal{Q}_{i,t}$. We call “nearest-query” this distance. We compare the performance of GHC with three distances: the nearest-query distance, the distance to Euclidean convex hulls, and the distance to spherical convex hulls.
@@ F.2.3 Additional Real-World Experiments
In Fig. 9, we present the performance of our models on the ComQA and CQADupStack datasets using the nearest-query distance. Nomic surprisingly outperforms Mistral_E5 on both datasets, achieving a lower accumulated regret of 1,870 compared to 2,081 on ComQA, and 7,549 compared to 7,586 on CQADupStack. We also observe that Mistral_E5 accumulates less regret than E5-large, which is an expected outcome given the design and capabilities of the two models.
@@
Figure 9: Comparison of GHC with different LLMs on ComQA and CQADupStack datasets.
@@ F.3 Compute Resources for Experiments
Throughout our experiments, the only component that required substantial computational resources was the encoding of our datasets using Mistral_E5. As Mistral_E5 is a Large Language Model comprising 7 billion parameters, it necessitates the use of a high-memory GPU for inference, such as the NVIDIA H100 with 80 GB of memory. The encoding of QQG and ComQA takes approximately 20 minutes while the encoding of CQADupStack takes approximately 4 hours. Every other component of the experiments can be executed on a single workstation, provided it is capable of running continuously for up to two days to accommodate the more computationally intensive tasks.

### v2

@@ F.2.1 Real-World Datasets
Quora Question Groups (QQG): Quora Question Groups is derived from the Quora Question Pairs dataset (Wang et al., 2017). The original dataset comprises over 400,000 question pairs, each annotated with a binary label indicating whether the questions are paraphrases of each other. To transform this pairwise paraphrase data into coherent question groups, we represented the dataset as a graph where nodes correspond to questions and edges indicate paraphrase relationships.
@@
ComQA [^6] https://paperswithcode.com/dataset/comqa (Abujabal et al., 2019): ComQA consists of 11,214 English questions collected from the WikiAnswers forum and grouped into 4,834 paraphrase clusters by crowd workers. By design, ComQA emphasizes reasoning-intensive phenomena—namely compositional queries, explicit temporal constraints, and comparative or superlative constructions—making it a rigorous benchmark for research on complex factoid question answering. The middle panel of Fig. 8 presents the group size distribution of this dataset.
@@
CQADupStack  [^7] https://github.com/D1Doris/CQADupStack (Hoogeveen et al., 2015): CQADupStack is a public benchmark available to researchers working on question-answering learning tasks. It consists of complete threads from twelve StackExchange subforums (android, english, gis, mathematica, physics, programmers, statistics, superuser, tex, unix, webmasters, wordpress) annotated with community‐marked duplicate links. To evaluate our classification approach across distinct expert domains, we apply it to the Mathematica, Physics, and Statistics parts of the dataset. The right panel of Fig. 8 presents the group size distribution of this dataset. The subset of the dataset we use comprises 99,785 questions organized into 74,519 groups. CQADupStack is the most challenging dataset used in our experiments, because of its relatively small average group size.
F.2.2 Text Embedding Models
@@
In this section, we describe the text embedding models selected for our experiments, along with the way they were trained and fine-tuned. We chose three models that are widely recognized in the literature and frequently cited due to their strong performance across a range of Natural Language Processing tasks.
@@
E5 (Wang et al., 2022): EmbEddings from bidirEctional Encoder rEpresentations (E5) is available in three versions. The first version is initialized with MiniLM (Wang et al., 2020), the second is initialized with BERT-base (Reimers and Gurevych, 2019), and the third is initialized with BERT-large. This model follows a bi-encoder architecture, where both the query and passage encoders are initialized with BERT. The training process consists of two stages. The first stage, called “unsupervised”, uses a large number of unlabeled data, including title and passage pairs from Wikipedia, questions and answers from Reddit, and more. The InfoNCE contrastive loss (van den Oord et al., 2018) is employed to minimize the distance between related queries and passages, while maximizing the distance between unrelated queries and passages. The second stage involves training the model on high-quality human-annotated data, such as NLI (Gao et al., 2021), MS-MARCO (Bajaj et al., 2016), and Natural Questions (Karpukhin et al., 2020). During this stage, the model is trained with a loss that combines the KL divergence between the probability distribution of the label, as given by a cross-encoder teacher model, and the probability distribution generated by our E5 student model, along with the InfoNCE contrastive loss. This second stage further improves model performance on benchmarks such as BEIR (Thakur et al., 2021) and MTEB (Muennighoff et al., 2022).
@@
NOMIC (Nussbaum et al., 2024): Nomic is initialized from BERT and modified to address long-context retrieval. Nomic consists of 100 million parameters and supports a sequence length of up to 2048. Nomic’s pretraining is divided into three stages. The first stage focuses on Masked Language Modeling to learn longer sequence representations. The subsequent stages are supervised and unsupervised, both employing the InfoNCE contrastive loss. This model was trained on a significantly larger dataset compared to the previous versions, encompassing both supervised and unsupervised phases. Nomic uses task-specific prefixes—such as search, search query document, classification, and clustering—to distinguish between the behaviors of each task. For the purpose of our work, we used the clustering prefix.
@@
Mistral_E5 (Wang et al., 2024): This is the first unidirectional decoder architecture we use for our work. The model is initialized from Mistral 7B Jiang et al. (2023) and consists of 7 billion parameters. The model takes as input the query $q^{+}$ and the task_definition and generates the instruction template:
@@
The query and document instruction templates are then passed to the model. We obtain the query and document embeddings, $h^{+}_{q}$ and $h^{+}_{d}$, from the last layer of the model at the [EOS] position. The model was trained on a large corpus of both original and synthetic data. The synthetic data was generated using advanced large language models such as GPT-4.
F.2.3 Additional Real-World Experiments on QQG
In this section, we evaluate the performance of GHC (Algorithm 2), CC and the Sequential $k$-Means (SKM) baseline using the datasets and embedding models described in Sections F.2.1 and F.2.2 respectively. The models we utilize are typically trained using a loss function based on cosine similarity. Since they output unit-normalized embeddings, maximizing cosine similarity is equivalent to minimizing Euclidean distance. Hence, the reliance of GHC on Euclidean distance naturally leverages the models’ cosine-trained embeddings. As explained in Section 4, we always initialize the algorithms by manually picking an example question from each group. For GHC, this skips the first phase where CHC is applied until every group is represented. We then run our algorithms on the remaining questions in the dataset considered.
To align with the Retrieval-Augmented Generation (RAG) literature (Lewis et al., 2020) and improve computational efficiency, it is sensible to consider the following modification of the GHC algorithm: instead of computing distances to the convex hulls $\hat{\mathcal{C}}_{i,t}$ of each set of expert queries $\mathcal{Q}_{i,t}$, we compute the distance to the nearest query in $\mathcal{Q}_{i,t}$. We refer to this version of GHC as “nearest-query GHC”. We compare the performance of GHC with three distances: the distance to the nearest query, the distance to Euclidean convex hulls, and the distance to spherical convex hulls.
@@ F.2.3 Additional Real-World Experiments on QQG
In Fig. 10, we present the performance of our models on the ComQA and CQADupStack datasets using the nearest-query version of GHC. Nomic surprisingly outperforms Mistral_E5 on both datasets, achieving a lower accumulated regret of 1,870 compared to 2,081 on ComQA, and 7,549 compared to 7,586 on CQADupStack. We also observe that Mistral_E5 accumulates less regret than E5-large, which is an expected outcome given the design and capabilities of the two models.
@@
Figure 10: Comparison of GHC with different text embedding models on ComQA and CQADupStack datasets.
F.2.4 Comparison of GHC with CC and baselines
In this section, we extend the comparative analysis of Section 4 and compare the nearest-query version of GHC (see Section F.2.3) with CC for $C=0$ (see Section 3.2), AMP and SKM  [^8] See Appendix F.1.1 for a description of these baselines. on the QQG and ComQA datasets, using all three embedding models. As in Section 4, we perform three independent runs, and only report the average cumulative regret of GHC and AMP for the best performing threshold among specific ranges, which are:
  1. (i)
    $S=\{0,0.10,0.20,0.40,0.60,0.70,0.80,0.85,0.90,0.95,0.99,1\}$ for GHC
  2. (ii)
    $S\cup\{0.92,0.94,0.96,0.98\}$ for AMP (because AMP exhibited better performance with very high thresholds)
The results, which are reported in Fig. 12, provide additional insight into the relative difficulty of the QQG and ComQA datasets across different algorithms and embedding models. Overall, ComQA appears to be a more challenging dataset, as evidenced by consistently higher regret for all algorithms and all models when compared to QQG. This trend holds across both Mistral and Nomic models, indicating that the increased difficulty is dataset-intrinsic rather than model-specific. When comparing embedding models, Mistral_E5 achieves the best overall performance on QQG, yielding lower accumulated regret across all algorithms. In contrast, Nomic performs best on the ComQA dataset, suggesting stronger robustness to the increased complexity of ComQA-style queries. In terms of algorithmic performance, the AMP baseline exhibits sublinear regret on QQG, while showing approximately linear regret on ComQA, highlighting its difficulty in adapting to the harder setting. This indicates that AMP struggles to effectively leverage feedback under increased ambiguity and noise. In contrast, our proposed algorithm GHC consistently achieves the lowest regret across all datasets and embedding models. Moreover, the performance gap between GHC and competing baselines increases over time, particularly on ComQA, demonstrating superior robustness in challenging environments.
(a) Regret vs. step, Mistral_E5, QQG.
(b) Regret vs. step, Mistral_E5, ComQA.
(c) Regret vs. step, Nomic, QQG.
(d) Regret vs. step, Nomic, ComQA.
(e) Regret vs. step, E5, QQG.
(f) Regret vs. step, E5, ComQA.
Figure 12: Comparison of GHC with baselines using different text embedding models on QQG and ComQA datasets.
@@ F.3 Compute Resources for Experiments
Throughout our experiments, the only component that required substantial computational resources was the encoding of our datasets using Mistral_E5. As Mistral_E5 is a text embedding model comprising 7 billion parameters, it necessitates the use of a high-memory GPU for inference, such as the NVIDIA H100 with 80 GB of memory. The encoding of QQG and ComQA takes approximately 20 minutes while the encoding of CQADupStack takes approximately 4 hours. Every other component of the experiments can be executed on a single workstation, provided it is capable of running continuously for up to two days to accommodate the more computationally intensive tasks.